In a sample of 49 adolescents who served as the subjects in an


immunologic study, one variable of


interest was the diameter of skin test reaction to an antigen. The sample mean


and standard deviation


were 21 and 11mm erythema, respectively. Can it be concluded from these


data that the population


mean is less than 30? Let alpha = 0.5

1
Expert's answer
2022-03-21T02:01:12-0400

"H_0:\mu \geqslant 30\\H_1:\mu <30\\T=\sqrt{n}\frac{\bar{x}-\mu}{s}=\sqrt{49}\frac{21-30}{11}=-5.72727\sim t_{n-1}=t_{48}\\P-value:\\P\left( T<-5.72727 \right) =F_{t,48}\left( -5.72727 \right) =3.26\cdot 10^{-7}"

Since the P-value is less than the significance level, the null hypothesis is rejected. The mean value is less than 30.